OpenStudy (josee):

give a rigorous proof that f(x)= x^2+3x is continous at x=2

OpenStudy (anonymous):

f(2) = 4+6=10 , lim f(x) [when x___> 2 ] = 4+6 =10 , ===> f(x) is continous at x=2

OpenStudy (josee):

when it says rigorous proof, you hae to show it by using epsilon and delta :P

OpenStudy (anonymous):

 there's only 1 point on the line when x=2 which makes it continuous right?